OpenStudy (anonymous):

Simplify the expression 11x − (2x^2 + 8x).

OpenStudy (here_to_help15):

lemme see

OpenStudy (anonymous):

1. Simplify brackets 11x−2x2−8x 2. Gather like terms (11x−8x)−2x2 3. Simplify 3x−2x2 <---- :)

OpenStudy (here_to_help15):

it is x(3-2x)

OpenStudy (anonymous):

@Here_to_Help15 wrong

OpenStudy (anonymous):

@Here_to_Help15 were are your steps

OpenStudy (here_to_help15):

no i am not!

OpenStudy (here_to_help15):

i'll show them

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

11x − (2x^2 + 8x) 11x-2x^2-8x (11x-8x) -2x^2 3x-2x^2 :)

OpenStudy (here_to_help15):

11x-(2x^2+8x) Factor 2x out of 2x^2 +8x 11x - 2x(x+4) Factor x out of 11x - 2x(x+4),resulting in x(11 - 2(x + 4) -2(x + 4)=-8-2x x(-8-2x+11) Grouping like terms,-2x -8+11=(11-8)-2x 11-8=3 Answer: x(3-2x)

OpenStudy (anonymous):

lol i just said it in the the thread before. @tayah.val.17 just foolow that and make sure you understand it! thats the key

OpenStudy (here_to_help15):

well geez :(

OpenStudy (anonymous):

;( did i sound to harsh?

OpenStudy (anonymous):

@Here_to_Help15 im realy sorry if I sounded mean, I had a very bad day. And we went to a funeral yesterday DX and @tayah.val.17 here to help is right. Listen to their's.

OpenStudy (anonymous):

no you didn't it helped me? thanks

OpenStudy (anonymous):

aww sorry to hear that happend

OpenStudy (anonymous):

no its ok its not your fault. He had some problem with him and he died in the hospital..

OpenStudy (anonymous):

ohh that sucks. my great grandma died at 92 in her bed. last year

OpenStudy (anonymous):

oh my! thats so sad. I can feel your pain.

OpenStudy (anonymous):

yea i can feel yours

OpenStudy (anonymous):

its so sad to talk about those things. lets not talk about them. i need sometime to get over it..

OpenStudy (anonymous):

Given the ordered pair (3, y), what value of y would make the ordered pair a solution of the equation 4x − 2y = 24?

OpenStudy (anonymous):

4x − 2y = 24 4(3) -2y=24 12-2y=24 2y=(24-12) 2y=12 y=12/2 y=6 I hope you get it right!

OpenStudy (anonymous):

and i hope im not wrong :[

OpenStudy (anonymous):

i got it thanks. i suck at algebra/ math

OpenStudy (anonymous):

lol i use to thnk that. but soon you will realize it may be hard but you survived! :D

OpenStudy (anonymous):

yea. i am going to 10 grade will it be harder

OpenStudy (anonymous):

im in geometry right now. and its sad to say, but they barely use any algebra!! so when i took algebra it felt like it wasnt even needed! but you still need it for algebra 2

OpenStudy (anonymous):

your older than me. >:[ no fair lol jkjk

OpenStudy (anonymous):

im going into 9th grade and im taking algebra 2 i think ;( im scard.

OpenStudy (anonymous):

oh. well i am 15 you

OpenStudy (anonymous):

13

OpenStudy (anonymous):

but im turning 14. but by the end of this year. so im not really suppose to be going into 9th. but i dont understadn the month thingie about after september

OpenStudy (anonymous):

so idc lol!

OpenStudy (anonymous):

( idc about the month thingie)